OpenStudy (anonymous):

write the polynomial y = x^3 - 36x in factored form?

OpenStudy (anonymous):

y = x ( x^2 - 36)

OpenStudy (anonymous):

a^2 - b^2 = (a+b) (a-b)

OpenStudy (anonymous):

x^2 - 36 = (x+6) ( x-6) y = x (x+6) (x-6)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

Welcome.:)